An insurance company offers its policyholders a number of different premium payment

options. For a randomly selected policyholder, let X = number of months between

successive payments. The cdf of is as follows:

0  x < 1

0.30 1 ≤ x< 3

f(x)= 0.40 3 ≤ x < 4

0. 45 4 ≤ x < 6

0.60 6 ≤ x ≤ 12

1 12 ≤ x .


a) What is the pmf of X ?

b) Using just the cdf, compute P(3 ≤ X ≤ 6) and P(4 ≤ X).


1
Expert's answer
2021-09-27T15:11:15-0400

Solution:

We are given with cumulative distribution function:


"F(x)-\left\{\begin{array}{cc}0 & x<1 \\ 0.3 & 1 \leq x<3 \\ 0.4 & 3 \leq x<4 \\ 0.45 & 4 \leq x<6 \\ 0.60 & 6 \leq x<12 \\ 1 & 12 \leq x\end{array}\right."

 (a)

At every step, by subtracting subsequent cumulative distribution function, we can find probability distribution function.

"p(0)=0
\\p(1)=F(1)-F(0)=0.3-0=0.3
\\p(2)=0
\\p(3)=F(3)-F(2)=0.4-0.3=0.1
\\p(4)=F(4)-F(3)=0.45-0.40=0.05
\\p(5)=0
\\p(6)=F(6)-F(5)=0.60-0.45=0.15
\\p(6)=0
\\p(7)=0
\\p(8)=0
\\p(9)=0
\\p(10)=0
\\p(11)=0
\\p(12)=F(12)-F(11)=1-0.60=0.40"

(b)

We have to find probability between x=3 and x= inclusive. This is given by:

"\begin{aligned}

&p(3 \leq x \leq 6)=F(6)-F(2) \\

&=0.60-0.30 \\

&=0.30

\end{aligned}"

Therefore, "p(3 \leq x \leq 6)=0.30"


Also, we have to find probability for "x" greater than or equal to 4 . This is given by:

"\begin{aligned}

&p(4 \leq x)=F(12)-F(3) \\

&=1-0.40 \\

&=0.60

\end{aligned}"

Therefore,

"p(4 \leq x)=0.60"